If i+2j+k & 3i+2j+k are two sides of the parallelogram, then magnitude of the area of the parallelogram?

Grade:12th Pass

1 Answers

anila
6 years ago
Area of the parallelogram=|a×b| i.e; do the determinant of the given two sides of parallelogram u will get the answer as 2j-4k then magnitude=√2^2+4^2=5√2